when x is subtracted from 2y, the difference is equal to the average of x and y What is the value of x/y?
1
Expert's answer
2012-08-14T07:33:39-0400
We can construct the following equation:
2y-x=(x+y)/2
<=> 4y-2x=x+y
<=> 3y=3x
<=> y=x
Therefore x/y=1